OpenStudy (anonymous):

Find all values of x for which the function is differentiable. cuberoot(3x-6)+5

OpenStudy (anonymous):

\[\sqrt[3]{3x-6}+5\]

OpenStudy (anonymous):

d/dx=(1/3)*(3x-6)^(-2/3)*3 =(3x-6)^(-2/3) then it will be differentiabel over x=R-{2}

OpenStudy (anonymous):

Huh? I get up to this part..\[(1/3)(3x-6)^{(-2/3)}\] So how do I find the values for which x is differentiable?

OpenStudy (anonymous):

you need to write what that thing really is

OpenStudy (anonymous):

it is \[\frac{1}{3\sqrt[3]{(3x-6)^2}}\]

OpenStudy (anonymous):

and since you have a denominator it cannot be zero, in other words \[3x-6\neq 0\] \[x\neq2\]

OpenStudy (anonymous):

I think I need to go back to algebra again... What happened to the square and the cuberoot?

OpenStudy (anonymous):

\[b^{\frac{p}{q}}=\sqrt[q]{b^p}\]

OpenStudy (anonymous):

Oh yes I understand that part.

OpenStudy (anonymous):

and the minus sign puts it in the denominator

OpenStudy (anonymous):

oh maybe you mean what happened to them when i solved?

OpenStudy (anonymous):

Ohh nevermind, I see now. ?